OpenStudy (anonymous):

a. Suppose a market research company finds that at a price of p = $20, they would sell x = 42 tiles each month. If they lower the price to p = $10, then more people would purchase the tile, and they can expect to sell x = 52 tiles in a month’s time. Find the equation of the line for the demand equation. Write your answer in the form p = mx + b. Hint: Write an equation using two points in the form (x,p).

OpenStudy (anonymous):

if we know that when p =$20, x=42 we can find: p = mx + b 20 = m(41) + b (1) and from p=$10 when x=52, we can find: p = mx + b 10 = m(52) + b (2) use eqt. (1) to find b: b = 20 - 42(m) apply b into eqt. (2): 10 = 52(m) + 20 - 42(m) 52(m) - 42(m) = 10 - 20 10(m) = -10 [ m = -1 ] ans. with m, form the eqt.: p = -x + b (3) plug either points (42,20) or (52,10) to find b: 20 = -42 + b b = 62 10 = -52 + b b = 62 hence, p = -x + 62 is that right?